If you're given that tan θ = 7/2 and θ is an acute angle, locate the 5 remaining trigonometric functions of θ. Thanks.
Answer by Cromlix(4381)

Hi there,
tanθ = 7/2
cosθ = 2/√53
sinθ = 7/√53
cotangentθ = 2/7
secantθ = √53/2
cosecantθ = √53/7
Hope this helps :-)